OpenStudy (tkbellapie):

Justin is married with one child. He works 40 hours each week at a rate of $16 per hour. His wife began working part time after their daughter was born, but still contributes about $350 to the cash inflow each month. Their monthly cash outflow is generally about $3,000. They have a balance of $2,000 in their savings account. Justin has retirement contributions taken out of his paycheck at work. They have renter’s, car and life insurance coverage. Based on this information, what part of their financial plan should Justin and his wife work on?

OpenStudy (retireed):

Their combined income per month is around $3125 pre 401K, federal tax, FICA and Medicare, so I don't know if the $3000 monthly outflow includes paying the 401K, the taxes and the insurance payments. Worse case, with only $2000 in savings, they are one disaster away from bankruptcy, so he needs them to reduce their monthly outflow and getting an emergency fund built up of six months worth of income, so they need to save up around $18750. If he can get them to have an outflow of $2500, banking the extra $500. They could have the emergency fund up to $18750 in 18750 - 2000 = 16750/500 = 33 months With the savings fund at a reasonable level, the extra cash per month could be put into increasing the 401K contributions. Hopefully, he was putting in the company matching value, say 4% and the company 4%. 40 *16 *4.333* 0.04 = 111 per month times 33 months = $3633 added to the 401K with his money plus $3633 of company money So now Justin and his wife have increased their worth by $24075 in less than 3 years Justin can now put the additional $500 per month towards retirement. Just kidding. That would make his retirement saving rate of around 22.5% plus the 4% company match, which would be way too much to save.